A man invests $12,000 in two accounts. One account pays 10% per year. The other account pays 7% per year. How much was invested in each account, if the total interest earned in the first year was $960
Answer by mananth(16949)

part I 10.00% per annum ---x
part II 7.00% per annum ---y

x+y=12000 ------------------------1
10.00% x + 7.00% y= 960
Multiply by 100
10x+7y= 96000 --------2

Multiply (1) by -10
we get

-10x-10y=-120000.00

Add this to (2)

-3y= = -24000

divide by -3

y= 8000 investment at 7.00%
Balance 4000 investment at 10.00%
CHECK
8000 --------- 7.00% ------- 560.00
4000 ------- 10.00% ------- 400.00
Total -------------- 960.00